OpenStudy (anonymous):

A rain gutter is to be constructed of aluminum sheets 12 inches wide. After marking off a length of 4 inches from each edge, this length is bent up at an angle Θ. The area A of the opening may be expressed as the function: A(Θ) = 16 sin Θ • (cos Θ + 1). If Θ = 60°, what is the area of the opening?

OpenStudy (anonymous):

here are the answers: 1.2 in2 7.9 in2 16.3 in2 20.8 in2

OpenStudy (anonymous):

@gabylovesu

OpenStudy (anonymous):

I am thinking D or C

OpenStudy (anonymous):

can you show me how

OpenStudy (anonymous):

Nevermind I am thinking B

OpenStudy (anonymous):

A(45°) = 16sin(45°)(cos(45°) + 1)    = 16(1/√2)(1 + 1/√2)    = 16(1/√2)(1 + √2)/√2    = 16(1 + √2)/2    = 8 + (√2)/2 in²

OpenStudy (anonymous):

that's not one of my answer choices

OpenStudy (anonymous):

@gabylovesu

OpenStudy (anonymous):

@jdoe0001

OpenStudy (anonymous):

@OOOPS

OpenStudy (anonymous):

@mathstudent55

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@triciaal

OpenStudy (anonymous):

@Catch.me

OpenStudy (anonymous):

need pic, can you draw it?

OpenStudy (anonymous):

thats all there is

OpenStudy (anonymous):

ok, no need anymore, just plug \(\theta =60\) into the function sin(60) = \(\sqrt3/2\) cos (60)=1/2

OpenStudy (anonymous):

that's still not one of the answers

OpenStudy (anonymous):

sure, I gave you the way to do, just plug in and hit the keys of calculator to get the answer. It's your duty, not me.

OpenStudy (anonymous):

do you know what he's talkin about @gabylovesu

OpenStudy (anonymous):

um not sure

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@blahhedup101

OpenStudy (anonymous):